OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x - y = 6 -4x - 4y = 10

OpenStudy (anonymous):

Multiply the second equation by -4. Add the two equations together. Multiply the first equation by 4. Multiply the first equation by -4.

OpenStudy (anonymous):

Multiply the second equations by -4

OpenStudy (anonymous):

You shouldn't multiply the second equation by -4 because it wouldn't cancel out the variable x. You should multiply the first equation by 4. Then you will have this: 4x - 4y = 24 -4x - 4y = 10 And the x variables will cancel each-other out and you can solve for y.